OpenStudy (moonlitfate):

Assuming that Earth is a sphere of radius 6378 kilometers, what is the difference in the latitudes of Syracuse, New York and Annapolis, Maryland, where Syracuse is about 450 kilometers due north of Annapolis?

OpenStudy (campbell_st):

you have the arc length and radius.... just need to find the value of theta \[l = r \theta \] where theta is the subtended angle...in radians..

OpenStudy (moonlitfate):

So, theta is -- \[\theta = s/r\] \[\theta = 6878/450\]

OpenStudy (moonlitfate):

*6378/450

OpenStudy (campbell_st):

other way round.... \[\theta = 450/6378\] that will give an answer in radians. To convert radians to degrees multiply by \[\frac{180}{\pi}\]

OpenStudy (campbell_st):

its around 4 degrees..

OpenStudy (moonlitfate):

Ah, I see where I messed up! Thank you. :)

OpenStudy (campbell_st):

good luck